OpenStudy (camzzzie):

HELP PLEASE? WILL MEDAL The half-life of a certain radioactive material is 42 days. An initial amount of the material has a mass of 49 kg. Write an exponential function that models the decay of this material. Find how much radioactive material remains after 8 days. Round your answer to the nearest thousandth.

OpenStudy (anonymous):

i can help.just wait:)

OpenStudy (camzzzie):

Ok thank you

OpenStudy (anonymous):

I can help hold on

OpenStudy (anonymous):

does that help:)

OpenStudy (anonymous):

The half-life of a certain radioactive material is 42 days. An initial amount of the material has a mass of 49 kg. Write an exponential function that models the decay of this material. Find how much radioactive material remains after 8 days. Round your answer to the nearest thousandth. y = 49(1/2)42x;0 kg y = 1/2(1/49)1/42x;0.238 kg y = 49(1/2)1/42x;42.940 kg y = 2(1/49)1/42x;0.953 kg Use the graph of y = ex to evaluate e1.6 to four decimal places. 4.3493 4.9530 0.2019 2.7183

OpenStudy (camzzzie):

these are the options given y = 0 kg y = ; 0.238 kg y = ; 42.940 kg y = ; 0.953 kg

OpenStudy (michele_laino):

I think that, if m(t) is the mass of your radioactive sample at time t, and m_0 is the mass of your radioactive sample at initial time, then we have: \[m(t)=m _{0}e ^{-t/\tau} \] where tau=42/ln2 m_0=49

OpenStudy (michele_laino):

I think m(t)=42.940 Kg, so the third option

OpenStudy (michele_laino):

sorry: m(8)=42.942 Kg @camzzzie

OpenStudy (camzzzie):

answer was c